Find the slope-intercept form for the line satisfying the following conditions. X-intercept 3, Y-intercept 2 over 3.
Answer by Cromlix(4381)

Slope = 'y' distance/'x' distance
= 2/3 / 3
= 2/9
Using the equation:
y = mx + c
Slope (m) = 2/9
c = 2/3
y = 2/9x + 2/3
OR multiply thro' by 9
9y = 2x + 6
Hope this helps.
:-)